# Learn With Imagination: Safe Set Guided State-wise Constrained Policy Optimization

### v4

@@ Abstract
Deep reinforcement learning (RL) excels in various control tasks, yet the absence of safety guarantees hampers its real-world applicability. In particular, explorations during learning usually results in safety violations, while the RL agent learns from those mistakes. On the other hand, safe control techniques ensure persistent safety satisfaction but demand strong priors on system dynamics, which is usually hard to obtain in practice. To address these problems, we present Safe Set Guided State-wise Constrained Policy Optimization (S-3PO), a pioneering algorithm generating state-wise safe optimal policies with zero training violations, i.e., learning without mistakes. S-3PO first employs a safety-oriented monitor with black-box dynamics to ensure safe exploration. It then enforces an "imaginary" cost for the RL agent to converge to optimal behaviors within safety constraints. S-3PO outperforms existing methods in high-dimensional robotics tasks, managing state-wise constraints with zero training violation. This innovation marks a significant stride towards real-world safe RL deployment.
@@ 1 Introduction
Reinforcement Learning (RL) has showcased remarkable advancements in domains like control and games. However, its focus on reward maximization sometimes neglects safety, potentially leading to catastrophic outcomes (Gu et al., 2022). To rectify this, the concept of safe RL emerged, aiming to ensure safety throughout or after training. Initial endeavors centered on Constrained Markov Decision Processes, often emphasizing cumulative or chance constraints (Ray et al., 2019; Achiam et al., 2017; Liu et al., 2021). While effective, these approaches lack instantaneous safety guarantees, which is crucial for managing emergencies like collision avoidance in autonomous vehicles (Zhao et al., 2023c; He et al., 2023a). Recent strides Zhao et al. (2023b) leveraged the Maximum Markov Decision Process to ensure instantaneous safety by bounding violations while integrating trust region techniques for policy enhancement, resulting in simultaneous improvement of worst-case performance and adherence to cost constraints. However, these approaches still cannot ensure safety during learning due to potentially unsafe explorative behaviors.
Meanwhile, safe control methods to continuously meet stringent safety requirements in predictable environments are widely examined, with energy function-based methods being the most prevalent Khatib (1986); Ames et al. (2014); Liu and Tomizuka (2014); Gracia et al. (2013); Wei and Liu (2019). These methods establish energy functions assigning lower energy to safe states and project nominal control into energy dissipating control, hence persistently maintaining system safety. However, these approahces heavily rely on accessible white-box analytical models of system dynamics. The Implicit Safe Set Algorithm (ISSA) Zhao et al. (2021, 2024) addresses this problem using black-box optimization over black-box dynamics models. Nevertheless, the reliance on continuous monitoring to ensure safety imposes significant computational overhead, and the RL policy itself does not inherently guarantee safe behavior. This combination presents a major obstacle to the successful deployment of the algorithm in real-time, safety-critical applications.
@@
In summary, a pressing need arises for a method leveraging both the zero-training-time-violation ability of safe control and optimal convergence of safe RL. To bridge this gap, we introduce a novel Safe Set Guided State-wise Constrained Policy Optimization (S-3PO) algorithm. S-3PO safeguards the exploration of immature policies with a black-box safe control method and derives a novel formulation where RL learns an optimal safe policy by constraining the state-wise “imaginary" safety violations. An overview of the prescribed S-3PO principles is presented in Figure 1. Empirical validation underscores S-3PO’s efficacy in training neural network policies encompassing thousands of parameters for high-dimensional simulated robot locomotion tasks. Our contribution marks a substantial advancement in the realm of practical safe RL algorithms, poised to find applications in a multitude of real-world challenges.
@@ 2.1 Preliminaries
To maintain simplicity, our approach focuses on deterministic dynamics, although it is worth noting that the proposed method can be readily extended to accommodate stochastic dynamics Zhao et al. (2021); Noren et al. (2021). Additionally, we assume the access to the dynamics model $f$ is only in the training phase, and is restricted to an implicit black-box form, as exemplified by an implicit digital twin simulator or a deep neural network model Zhao et al. (2021). We also assume there is no model mismatch, while model mismatch can be addressed by robust safe control Wei et al. (2022) and is left for future work. Post training, the knowledge of the dynamics model is concealed, aligning with practical scenarios where digital twins of real-world environments are too costly to access during model deployment.
@@
In this research, our primary focus lies in ensuring safety for episodic tasks, which falls within the purview of finite-horizon Markov Decision Processes (MDP). An MDP is defined by a tuple $(\mathcal{S},\mathcal{A},\gamma,R,P,\mu)$. The reward function is denoted by $R:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$, the discount factor by $10\leq\gamma<$ 1, the initial state distribution by $\mu:\mathcal{S}\mapsto\mathbb{R}$, and the transition probability function by $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathbb{R}$.
@@
The safety specification requires that the system state remains within a closed subset in the state space, denoted as the “safe set” $\mathcal{S}_{S}$. This safe set is defined by the zero-sublevel set of a continuous and piecewise smooth function $\phi_{0}:\mathbb{R}^{n_{s}}\rightarrow\mathbb{R}$, where $\mathcal{S}_{S}=\{s\mid\phi_{0}(s)\leq 0\}$. Users directly specify both $\mathcal{S}_{S}$ and $\phi_{0}$, which is easy to specify. For instance, for collision avoidance, $\phi_{0}$ can be specified as the negative closest distance between the robot and environmental obstacles.
@@ 2.2 Problem
We are interested in the safety imperative of averting collisions for mobile robots navigating 2D planes. We aim to persistently satisfy safety specifications at every time step while solving MDP, following the intuition of State-wise Constrained Markov Decision Process (SCMDP) Zhao et al. (2023c). Formally, the set of feasible stationary policies for SCMDP is defined as
@@
3 Prior Works
3.1 Safe Reinforcement Learning
Existing safe RL approaches either consider safety after convergence or safety during training Zhao et al. (2023c). End-to-end approaches are usually used to ensure safety after convergence Liang et al. (2018); Tessler et al. (2018); Bohez et al. (2019); Ma et al. (2021); He et al. (2023b). However, these approaches cannot avoid unsafe explorations. Safety in training is achieved by hierarchical approaches which uses a safeguard to filter out unsafe explorative actions. The safeguard relies on the knowledge on the system dynamics, which can be either learned dynamics Dalal et al. (2018); Thananjeyan et al. (2021); Zhao et al. (2022), white-box dynamics Fisac et al. (2018); Shao et al. (2021), or black-box dynamics Zhao et al. (2021). Nevertheless, these methods cannot guarantee convergence of the RL policy.
The proposed approach is the first to address safety both during training and after convergence, which can potentially serve as a general framework to bridge these two types of approaches. In the following, we choose the most advanced method in each category to form the proposed approach, so that it relies on the least assumptions.
@@ 3.2 Implicit Safe Set Algorithm
The Implicit Safe Set Algorithm (ISSA) Zhao et al. (2021, 2024) ensures the persistent satisfaction of safety specifications for systems with black-box dynamics (e.g., digital twins or neural networks) through energy function-based optimization. Leveraging energy function $\phi=\phi_{0}^{*}+k_{1}\dot{\phi}_{0}+\cdots+k_{n}\phi_{0}^{(n)}$ and theoretical results from SSA Liu and Tomizuka (2014), ISSA synthesizes a safety index to guarantee that the safe control set $\mathcal{A}_{S}(s):=\{a\in\mathcal{A}\mid\dot{\phi}\leq-\eta(\phi)\}$ is nonempty. And then the set $\bar{\mathcal{S}}:=\{s\mid\phi(s)\leq 0\}\cap\{s\mid\phi_{0}(s)\leq 0\}$ is forward invariant under mild assumptions. Here $\eta(\phi)$ is designed to be a positive constant when $\phi\geq 0$ and $\infty$- ∞ when $\phi<0$. Intuitively, the energy function $\phi$ ensures that low-energy states are safe, while the control signal within $\mathcal{A}_{S}(s)$ reduces $\phi$ whenever the system state is outside the safe region. This further derives the forward invariance of $\bar{\mathcal{S}}$, meaning that once the system enters the safe region, its state will stays within the safe part of the state space.
To match the notion of MDP, we consider discrete-time safe control set $\mathcal{A}_{S}^{D}(s):=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq\max\{\phi(s)-\eta,0\}\}$. Thus ISSA ensures safety by projecting the nominal control $a_{t}^{r}$ generated from RL policy $\pi_{\theta}$ to $\mathcal{A}_{S}^{D}(s_{t})$ through solving the following optimization:
$\begin{split}&\min_{a_{t}\in\mathcal{A}}\|a_{t}-a^{r}_{t}\|^{2}\\ &\text{s.t. }\phi(f(s_{t},a_{t}))\leq\max\{\phi(s_{t})-\eta,0\}.\end{split}$ (5)
The detailed mild assumoptions and safety index synthesis are summarized in Appendix A. And the algorithms for solving (5) are provided in Algorithms 2 and 3.
@@ 3.3 State-wise Constrained Policy Optimization
Safe RL algorithms under the framework of Constrained Markov Decision Process (CMDP) do not consider state-wise constraints. To address this gap, State-wise Constrained Policy Optimization (SCPO) was proposed Zhao et al. (2023b) to provide guarantees for state-wise constraint satisfaction in expectation, which is under the framework of State-wise CMDP (SCMDP). To achieve this, SCPO directly constrain the expected maximum state-wise cost along the trajectory. And they introduced Maximum MDP (MMDP). In this setup, a running maximum cost value is associated with each state, and a non-discounted finite MDP is utilized to track and accumulate non-negative increments in cost. The format of MMDP will be introduced in Section 4.
@@ 4 Safety Index Guided State-wise Constrained Policy Optimization
The core idea of S-3PO is to enforce zero safety violation during training by projecting unsafe actions to the safe set, and then constrain the "imaginary" safety violation (i.e., what if the projection is not done) to ensure convergence of the policy to an optimal safe policy.
@@
To ensure zero violation exploration, we safeguard nominal control via solving (5) at every time step during policy training. In this paper, we adopt ISSA to solve (5).
For the 2D collision avoidance problem considered in this paper, we choose the same safety index synthesis rule as [Section 4.1, (Zhao et al., 2021)], which is summarized in Appendix A. We show in Section 5 that with the safety index synthesis rule, ISSA is guaranteed to find a feasible solution of (5), making the system forward invariance in the set $\mathcal{\bar{S}}$.
@@
While eliminating safety violations during training is good, this also brings inevitable challenges for the policy learning as it directly eliminates the unsafe experience, making the policy unable to distinguish between safe and unsafe actions. To overcome this challenge, the main intuition we have is that instead of directly experiencing unsafe states, i.e. $s\notin\mathcal{S}_{S}$, policy can learn to act safely from “imagination", i.e., how unsafe it will be if the safeguard had not been triggered? The critical observation we rely upon is that:
Observation 1
Define $\Delta\phi_{t}=\Delta\phi(s_{t},a_{t},s_{t+1})\doteq\phi(f(s_{t},a^{r}_{t}))-\phi(f(s_{t},a_{t}))$, i.e. the degree of required correction to safeguard $a_{t}^{r}$. Therefore, $\Delta\phi_{t}$ can be treated as an imagination on how unsafe the reference action would be, where $\Delta\phi_{t}\leq 0$ means $a_{t}^{r}\in\mathcal{A}_{S}^{D}(s_{t})$.
Following 1, Equation 4 can be translated to:
@@
Policies satisfying (6) ensure there is no imaginary safety violation for any possible $a_{t}^{r}$, making $\pi_{\theta}$ a safe policy as required by (4).
@@ 4.1 Transfrom State-wise Constraint into Maximum Constraint
For (6), each state-action transition pair corresponds to a constraint, which is intractable to solve. Inspired by Zhao et al. (2023c), we constrain the expected maximum state-wise $\Delta\phi$ along the trajectory instead of individual state-action transition $\Delta\phi$.
Next, by treating $\Delta\phi_{t}$ as an “imaginary” cost, we define a MMDP Zhao et al. (2023c) by introducing (i) an up-to-now maximum state-wise cost $M$ within $\mathcal{M}\subset\mathbb{R}$, and (ii) a "cost increment" function $D$, where $D:(\mathcal{S},\mathcal{M})\times\mathcal{A}\times\mathcal{S}\mapsto[0,\mathbb{R}^{+}]$ maps the augmented state-action transition tuple to non-negative cost increments. We define the augmented state $\hat{s}=(s,M)\in(\mathcal{S},\mathcal{M})\doteq\hat{{\mathcal{S}}}$, where $\hat{{\mathcal{S}}}$ is the augmented state space. Formally,
@@
By setting $D\big{(}\hat{s}_{0},a_{0},\hat{s}_{1}\big{)}=\Delta\phi(s_{0},a_{0},s_{1})$, we have $M=\sum_{k=0}^{t-1}D\big{(}\hat{s}_{k},a_{k},\hat{s}_{k+1}\big{)}$ for $t\geq 1$. Hence, we define expected maximum state-wise cost (or $D$-return) for $\pi$:

### v5

@@ Abstract
Deep reinforcement learning (RL) has achieved remarkable success across various control tasks. However, its reliance on exploration through trial and error often results in safety violations during training. To mitigate this, safety filters are commonly employed to correct unsafe actions generated by the RL policy. Yet, a key challenge remains: how to enable safety-filter-guided learning to produce a policy that remains optimally safe even after the filter is removed. In this paper, we propose Safe Set Guided State-wise Constrained Policy Optimization (S-3PO) — a novel algorithm designed to generate optimal safe RL policies with zero training violations while maintaining safety during evaluation even without any safety filter. S-3PO integrates a safety-oriented monitor operating on black-box dynamics to ensure safe exploration and introduces an imaginary cost mechanism that guides the safe RL agent toward optimal behavior under safety constraints. This imaginary cost inherits the interpretability of the safety filter while outperforming conventional imitation-based cost designs. Equipped with state-of-the-art components, S-3PO demonstrates superior performance on high-dimensional robotic control tasks, effectively handling expected state-wise constraints and ensuring safety throughout the training process.
@@ 1 Introduction
Safe Reinforcement Learning (safe RL) has emerged as a powerful approach in domains such as games and robotic control, where ensuring safety during or after training is critical. While objective-based methods aim to optimize reward, they often lack formal guarantees on safety performance [Bohez et al. (2019)]. To address this, many approaches enforce hard constraints [Bouvier et al. (2024b), Bouvier et al. (2024a)]; however, these methods are typically effective only in low-dimensional systems. More recent advances [Zhao et al. (2023b), Zhao et al. (2024b)] leverage trust-region techniques combined with Maximum Markov Decision Processes (MMDP) to enforce simultaneous improvement of worst-case performance and adherence to cost constraints.
Despite these developments, RL-based methods fundamentally depend on trial-and-error exploration, making it difficult to guarantee safety throughout the training process. A common strategy to mitigate this issue is the use of safety filters [Alshiekh et al. (2018)], which is designed to correct unsafe actions generated by the RL policy. These safety filters are often constructed using principles from safe control theories [Shao et al. (2021)], where energy function-based methods remain the most widely adopted approach [Khatib (1986), Ames et al. (2014), Liu and Tomizuka (2014), Gracia et al. (2013), Wei and Liu (2019)]. Although safety filters can ensure safety during training by overriding unsafe actions, they also prevent the policy from learning how to avoid unsafe behaviors on its own. This creates a fundamental dilemma: how can an agent learn to avoid unsafe scenarios if it is always shielded from experiencing them?
To enable the policy to learn from the safety filter and generate safe actions by itself, rather than only being protected by the safety filter,  [Cheng et al. (2019)] use Gaussian Process (GP) models to estimate unknown system dynamics and construct a safety filter that implicitly guides the policy updates based on the history of safety filter interventions. To explicitly imitate the effect of the safety filter at each step, other works have considered using reward penalties to learn from safety filters.  [Krasowski et al. (2022)] introduce a constant penalty when the safety filter is activated. Yet, this approach does not account for how unsafe the proposed action was. In contrast,  [Wabersich and Zeilinger (2021)] penalize the reward with magnitude of the action correction applied by the safety filter. However, the magnitude of the correction alone may not accurately capture the true impact of the action on system safety, and the reward penalty can only impose a soft constraint during the learning process.
@@
These limitations highlight a critical gap: to enable a policy to maintain safety after the safety filter is removed, it is essential to introduce a cost term that explicitly measures how the policy’s actions influence the system’s safety level if the safe filter is removed. To this end, we introduce Safe Set Guided State-wise Constrained Policy Optimization (S-3PO). S-3PO safeguards the exploration of immature policies through a black-box safe control mechanism and formulates a novel constrained optimization framework where RL learns an optimal safe policy by constraining imaginary safety violations—violations that would have occurred without the filter.
@@ 2.1 Assumptions
To maintain simplicity, our approach focuses on deterministic dynamics, although the proposed method can be easily extended to stochastic dynamics [Zhao et al. (2021)]. Additionally, we assume the access to the dynamics model $f$ is only in the training phase and restricted to an black-box form, such as an implicit digital twin simulator or a deep neural network model [Zhao et al. (2021)]. We also assume there is no model mismatch. Model mismatch can be addressed by robust safe control [Wei et al. (2022)] and is left for future work. Post training, the knowledge of the dynamics model is concealed—a benefit of using "imaginary cost"—aligning with practical scenarios where digital twins of real-world environments are too costly to access during deployment. Based on these, our core target is to figure out how can safety-filter-guided learning be used to produce a policy that remains optimally safe even without the filter.
@@
In this research, our primary focus lies in ensuring safety for episodic tasks, which falls within the purview of finite-horizon Markov Decision Processes (MDP). An MDP is defined by a tuple $(\mathcal{S},\mathcal{A},\gamma,R,P,\mu)$. The reward function is denoted by $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$, the discount factor by $10\leq\gamma<$ 1, the initial state distribution by $\mu:\mathcal{S}\rightarrow\mathbb{R}$, and the transition probability function by $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$.
@@
The safety specification requires that the system state remains within a closed subset in the state space, denoted as the “safe set” $\mathcal{S}_{S}$. This safe set is defined by the zero-sublevel set of a continuous and piecewise smooth function $\phi_{0}:\mathbb{R}^{n_{s}}\rightarrow\mathbb{R}$, where $\mathcal{S}_{S}=\{s\mid\phi_{0}(s)\leq 0\}$, usually specified by users. For instance, for collision avoidance, $\phi_{0}$ can be specified as $d_{min}-d$ where $d$ is the closest distance between the robot and environmental obstacles and $d_{min}$ is the distance margin.
@@ 2.2 Problem
We are interested in the safety imperative of averting collisions for mobile robots navigating 2D planes. We aim to persistently satisfy safety specifications at every time step while solving MDP, following the intuition of State-wise Constrained Markov Decision Process (SCMDP) [Zhao et al. (2023c)]. Formally, the set of feasible stationary policies for SCMDP is defined as
@@
3 Preliminary
@@ 3.1 Implicit Safe Set Algorithm
As a deterministic safety filter, Implicit Safe Set Algorithm (ISSA) [Zhao et al. (2021), Zhao et al. (2024a)] ensures the persistent satisfaction of safety specifications for systems with black-box dynamics (e.g., digital twins or neural networks) through energy function-based optimization. Leveraging energy function $\phi=\phi_{0}^{*}+k_{1}\dot{\phi}_{0}+\cdots+k_{n}\phi_{0}^{(n)}$ and theoretical results from SSA [Liu and Tomizuka (2014)], ISSA synthesizes a safety index to guarantee that the safe control set $\mathcal{A}_{S}(s):=\{a\in\mathcal{A}\mid\dot{\phi}\leq-\eta(\phi)\}$ is nonempty. And then the set $\bar{\mathcal{S}}:=\{s\mid\phi(s)\leq 0\}\cap\{s\mid\phi_{0}(s)\leq 0\}$ is forward invariant under Assumption 1. Here $\eta(\phi)$ is designed to be a positive constant when $\phi\geq 0$ and $\infty$- ∞ when $\phi<0$.
Assumption 1
The system (1) is a second-order mobile platform that avoids 2D obstacles. 1) The state space is bounded, and the relative acceleration $w$ and angular velocity $z$ to the obstacle are bounded and both can achieve zeros, i.e., $w\in[w_{min},w_{max}]$ for $w_{min}\leq 0\leq w_{max}$ and $z\in[z_{min},z_{max}]$ for $z_{min}\leq 0\leq z_{max}$; 2) For all possible values of $z$ and $w$, there always exists a control $a$ to realize such $z$ and $w$; 3) The discrete-time system time step $dt\to 0$; 4) At any given time, there can at most be one obstacle becoming safety critical (Sparse Obstacle Environment).
Remark 3.1.
The bounds in the first assumption will be directly used to synthesize $\phi$. The second assumption enables us to turn the question on whether there exists a feasible control in $\mathcal{A}_{S}^{D}$ to the question on whether there exists $z$ and $w$ to decrease $\phi$. The third assumption ensures that the discrete time approximation error is small. The last assumption ensures that the safety index design rule is applicable to multiple moving obstacles.
The deterministic ISSA could be used as a safety filter in the discrete-time MDP and is treated as part of the environment during training. We define the discrete-time safe control set as $\mathcal{A}_{S}^{D}(s):=\{a\in\mathcal{A}\mid\phi(f(s,a))\leq\max\{\phi(s)-\eta,0\}\}$. The ISSA mechanism ensures safety by projecting the nominal control action $a_{t}$, proposed by the RL policy $\pi_{\theta}$, onto the safe control set $\mathcal{A}_{S}^{D}(s_{t})$ by solving the optimization problem:
$\begin{split}&\min_{a_{t}^{*}\in\mathcal{A}}\|a_{t}^{*}-a_{t}\|^{2}\\ &\text{s.t. }\phi(f(s_{t},a_{t}))\leq\max\{\phi(s_{t})-\eta,0\}.\end{split}$ (5)
@@ 3.2 State-wise Constrained Policy Optimization
Safe RL algorithms under the framework of Constrained Markov Decision Process (CMDP) do not consider state-wise constraints. To address this gap, State-wise Constrained Policy Optimization (SCPO) was proposed [Zhao et al. (2023b)] to provide guarantees for state-wise constraint satisfaction in expectation, which is under the framework of State-wise CMDP (SCMDP). To achieve this, SCPO directly constrain the expected maximum state-wise cost along the trajectory. And they introduced Maximum MDP (MMDP). In this setup, a running maximum cost value is associated with each state, and a non-discounted finite MDP is utilized to track and accumulate non-negative increments in cost. The format of MMDP will be introduced in Section 4.
@@ 4 Safety Index Guided State-wise Constrained Policy Optimization
Figure 2: (a) S-3PO pipeline during training. (b) S-3PO pipeline during evaluation.
The core idea of S-3PO is to enforce zero safety violations during training by projecting unsafe actions to the safe set, and then constraining the "imaginary" safety violations to ensure convergence of the policy to an optimal safe policy. As shown in Figure 2, the safety filter is part of the training environment, while the evaluation environment does not have the safety filter.
4.1 Learn with Imaginary Cost
@@
To ensure zero violation exploration, we adopt ISSA as the safety filter and safeguard nominal control via solving (5) at every time step during policy training. With the safety index synthesis rule in [Zhao et al. (2021)], ISSA is guaranteed to find a feasible solution of (5), making the system forward invariance in the set $\mathcal{\bar{S}}$. It is worth mentioning that any energy-function-based method that ensures forward invariance could be used as the safety filter, and the scalar energy function could be used to evaluate the imaginary cost. The integration with other energy-function-based methods will be left for future work.
@@
While eliminating safety violations during training is beneficial, it also poses challenges for RL policy training, as RL relies on a trial-and-error process. To address this, our key insight is that instead of directly encountering unsafe states ($s\notin\mathcal{S}_{S}$), the policy can leverage an "imaginary cost" to learn about unsafe scenarios without actually experiencing them.
Imaginary Cost
Define "imaginary cost" as $\Delta\phi_{t}=\Delta\phi(s_{t},a_{t},s_{t+1})\doteq\phi(f(s_{t},a_{t}))-\phi(f(s_{t},a_{t}^{*}))$, i.e. the degree of required correction to safeguard $a_{t}$. Here $a_{t}^{*}$ is the projected action by ISSA. Therefore, $\Delta\phi_{t}$ can be treated as an imagination on how unsafe the reference action would be, where $\Delta\phi_{t}\leq 0$ means $a_{t}\in\mathcal{A}_{S}^{D}(s_{t})$.
Following the definition, Equation 4 can be translated to:
@@
Policies satisfying (6) ensure there is no imaginary safety violation in expectation for any possible $a_{t}$, making $\pi_{\theta}$ a safe policy as required by (4), to be proved by lemma 5.1.
@@ 4.2 Transfrom State-wise Constraint into Maximum Constraint
For (6), each state-action transition pair corresponds to a constraint, which is intractable to solve. Inspired by [Zhao et al. (2023c)], we constrain the expected maximum state-wise $\Delta\phi$ along the trajectory instead of individual state-action transition $\Delta\phi$.
Next, by treating $\Delta\phi_{t}$ as an “imaginary” cost, we define a MMDP [Zhao et al. (2023c)] by introducing (i) an up-to-now maximum state-wise cost $M$ within $\mathcal{M}\subset\mathbb{R}$, and (ii) a "cost increment" function $D$, where $D:(\mathcal{S},\mathcal{M})\times\mathcal{A}\times(\mathcal{S},\mathcal{M})\rightarrow[0,\mathbb{R}^{+}]$ maps the augmented state-action transition tuple to non-negative cost increments. We define the augmented state $\hat{s}=(s,M)\in(\mathcal{S},\mathcal{M})\doteq\hat{{\mathcal{S}}}$, where $\hat{{\mathcal{S}}}$ is the augmented state space. Formally,
@@
By setting $D\big{(}\hat{s}_{0},a_{0},\hat{s}_{1}\big{)}=\Delta\phi(s_{0},a_{0},s_{1})$, we have $M=\sum_{k=0}^{t-1}D\big{(}\hat{s}_{k},a_{k},\hat{s}_{k+1}\big{)}$ for $t\geq 1$. Hence, we define expected maximum state-wise cost (or $D$-return) for S-3PO policy $\pi$: